OpenStudy (anonymous):

the lengths of the sides of a triangle are x, 3 and 5 what is the possible vaule of x?

OpenStudy (anonymous):

4

OpenStudy (anonymous):

why

OpenStudy (anonymous):

triples. :l

OpenStudy (anonymous):

4 if it is a right angled triangle

OpenStudy (anonymous):

yea abhed is right....theres a theorem where it says somthing about the addition or subtraction of the sides or somthing has to be equal to a certain quantity

OpenStudy (anonymous):

x^2+3^2=5^2 x^2=25-9 x^2=16 x=4

OpenStudy (anonymous):

ok so if a triangles sisdes are x, 33, and 41 how would i solve this one?

OpenStudy (anonymous):

pythagorean theorem. Should've remembered that. is x the hypotneuse? or however you spell that.

OpenStudy (anonymous):

no because the triangle is not drawn

OpenStudy (anonymous):

x can be taken as the hypotenuse only if it is th largest side

OpenStudy (anonymous):

If it's unspecified, how would you solve that?

OpenStudy (anonymous):

add 33+41?

OpenStudy (anonymous):

Assuming X is the hypotenuse, you would do: \[33^{2}+41^{2}=X ^{2}\]

OpenStudy (anonymous):

So X would be the square root of 2770, or 52.63078... There's probably a much simpler way to do that. But regardless, the answer is the same.

OpenStudy (anonymous):

there is.....all you need to do is add the two shortest side and they have to be greater than the largest side

OpenStudy (anonymous):

That's true, but how would it get you the answer? That would just mean X<74. Doesn't help much.